the probability that a trap will catch a rodent is 0.4 a) what is the probability that out of 5 traps set within a certain forest at least 1 will catch a rodent b) all will catch a rodent
Answer by stanbon(75887)

the probability that a trap will catch a rodent is 0.4
-------------------------------
a) what is the probability that out of 5 traps set within a certain forest at least 1 will catch a rodent
P(at least one) = 1 - P(none)
-------
= 1 - (0.6)^5
-----
= 1 - 0.0778
-------------------------
= 0.1846
--------------------------
b) all will catch a rodent
P(x = 5) = 0.4^5 = 0.01024
--------------------------------
Cheers,
Stan H.
===============================
-----------